# LPLC: A Dataset for License Plate Legibility Classification

### v1

@@ Abstract
Automatic License Plate Recognition (ALPR) faces a major challenge when dealing with illegible license plates (LPs). While reconstruction methods such as super-resolution (SR) have emerged, the core issue of recognizing these low-quality LPs remains unresolved. To optimize model performance and computational efficiency, image pre-processing should be applied selectively to cases that require enhanced legibility. To support research in this area, we introduce a novel dataset comprising 10,210 images of vehicles with 12,687 annotated LPs for legibility classification (the LPLC dataset). The images span a wide range of vehicle types, lighting conditions, and camera/image quality levels. We adopt a fine-grained annotation strategy that includes vehicle- and LP-level occlusions, four legibility categories (perfect, good, poor, and illegible), and character labels for three categories (excluding illegible LPs). As a benchmark, we propose a classification task using three image recognition networks to determine whether an LP image is good enough, requires super-resolution, or is completely unrecoverable. The overall F1 score, which remained below 80% for all three baseline models (ViT, ResNet, and YOLO), together with the analyses of SR and LP recognition methods, highlights the difficulty of the task and reinforces the need for further research. The proposed dataset is publicly available at https://github.com/lmlwojcik/lplc-dataset .
@@ I Introduction
Automatic License Plate Recognition (ALPR)plays a vital role in road surveillance by using automated systems to detect vehicles and recognize them based on their license plates (LPs) [1, 2]. In the deep learning era, key research tasks in this domain include LP detection and LP recognition, often complemented by additional processes such as vehicle detection and LP rectification [3, 4].
@@
State-of-the-art ALPR methods achieve over 95% accuracy in detection tasks and exceed 90% recognition rates under ideal conditions [5, 6]. However, their performance degrades significantly in challenging scenarios that are not well represented in most mainstream datasets. Such scenarios include low-light environments, adverse weather conditions (e.g., rain), and low-resolution or low-quality images caused by poor equipment or transmission compression artifacts [7, 8].
@@
Furthermore, recent advances in super-resolution (SR) technology have also opened up new paths to deal with faulty images [9, 10]. These models were adapted to an LP reconstruction task and were shown to aid in recovering information from otherwise illegible images. However, these models are often costly and not always needed, potentially even ruining otherwise suitable images for LP recognition (as we have observed in some of our experiments reported in this work).
@@
In this context, it is important to distinguish between image quality and character legibility. LP recognition performance depends not merely on image resolution or overall visual quality, but primarily on how legible the characters are. Fig. 1 illustrates this point: a visually high-quality image may contain illegible LPs due to factors such as camera distance, while a low-quality image may still feature clearly legible LPs. In some cases, a single image can include both legible and illegible LPs. We also observed that Optical Character Recognition (OCR) models often produce incorrect predictions with high confidence [11], particularly in cases where most characters are legible but one or two remain unclear (an issue also depicted in Fig. 1).
@@
With these challenges in mind, we present the License Plate Legibility Classification (LPLC) dataset111 The LPLC dataset is available at https://github.com/lmlwojcik/lplc-dataset , which contains over 10k radar images collected from various locations across the Brazilian state of Paraná and includes more than 12k annotated LPs. The images exhibit a wide range of capture conditions, including varying lighting scenarios, vehicle types (motorcycles, cars, buses, and trucks), and image quality. A key contribution of this dataset is the assignment of each LP to one of four legibility levels: perfect, good, poor, or illegible. These labels provide a qualitative and inherently subjective assessment of LP legibility. For the first three categories, we also provide OCR annotations. The dataset includes both Brazilian and Mercosur LPs 222Following prior literature [2, 11, 9], we use the term “Brazilian” to refer to the LP layout used in Brazil prior to the adoption of the Mercosur layout. and can be used as a benchmark for various ALPR-related tasks.
@@
In this work, we also use LPLC for a series of LP legibility assessment tasks using established image classification models: ResNet [12], ViT [13] and YOLO-cls [14]. Additionally, we explore the use of SR by evaluating three state-of-the-art models [9, 15, 10] on our novel dataset. Building on these components, we design a recognition pipeline that incorporates a new decision step to determine whether an input image is sufficiently legible, requires further processing, or should be considered unrecoverable.
@@ II Related Work
State-of-the-art advancements in ALPR, as in other machine learning research domains, are quantitatively assessed using public benchmark datasets [5, 3]. An example is CLPD [16], a dataset containing 1,200 images captured with various devices in mainland China. While it primarily features daylight scenes and passenger cars, it offers considerable variability in backgrounds, road types, capture devices, and image quality.
@@
For the Brazilian context, the RodoSol-ALPR dataset [5] comprises images captured at toll plazas, providing a realistic representation of actual operating conditions. It includes a wide range of lighting scenarios, vehicle types, and weather conditions. The authors [5] demonstrated that state-of-the-art methods perform well on older datasets featuring simpler scenarios, such as Caltech-Cars [17], but struggle on RodoSol-ALPR, revealing limitations tied to outdated benchmarks. While RodoSol-ALPR captures real situations, it is important to note that these toll plaza environments differ significantly from many surveillance camera settings, where camera quality and capture conditions are often much poorer and more variable.
@@
Recent advances in LP recognition have shifted focus from achieving high performance on simpler datasets to addressing more specific challenges in the field. For example, Liu et al. [6] proposed an attention-based decoder for LP recognition, following the growing trend of integrating attention mechanisms. Their approach combines a CNN encoder for feature extraction with a contrastive learning strategy designed to differentiate both the position and class of each character. The use of attention is motivated by the fixed character positions in LPs of a given format, enabling more effective layout-aware feature learning. They reported a % recognition rate on RodoSol-ALPR.
@@
Similarly, Rao et al. [18] proposed an end-to-end pipeline that includes a segmentation step. The method begins with a YOLOv5-based LP detector, followed by segmentation using the proposed AFF-Net model. Next, a skew correction module is applied to the segmentation map using standard image processing techniques to extract the LP region. The four corners of this region are then aligned with those of a canonical LP format. Finally, LP recognition is performed using a CNN. This correction step is effective in handling severely skewed LPs captured from oblique angles, such as those found in the CLPD dataset, where they reported a recognition rate of .
@@
SR has gained interest in ALPR research due to its potential to recover otherwise illegible LP images. For example, studies like [10] adapt image-to-image translation methods, such as Generative Adversarial Network (GAN), to improve recognition performance on low-resolution LPs.
@@
In the same direction, Nascimento et al. [9] proposed a novel focal loss tailored for character reconstruction, trained on a variant of the RodoSol-ALPR dataset called RodoSol-SR. This dataset comprises pairs of high-resolution LP images and their synthetically degraded low-resolution counterparts. These pairs were generated by the authors to train the model to reverse the degradation process through reconstruction.
@@ III The LPLC Dataset
We introduce the License Plate Legibility Classification (LPLC) dataset, a publicly available resource primarily designed for the task of classifying the legibility of LP images, that is, whether an LP image is suitable for direct OCR or requires additional processing. Although its main purpose is legibility classification, the dataset is also suitable for other tasks such as LP detection and recognition, due to its fine-grained annotations that include details on LP text, legibility levels, and occlusion. All annotations were made manually by a single person (the first author) and validated through a semi-automatic process to ensure accuracy. This revision process involved using an OCR model’s outputs (PARSeq-tiny [19]) and identifying legibility classification errors to flag potentially inconsistent annotations in both the legibility labels and OCR transcripts. Fig. 2 presents a few sample images from the dataset.
@@
The annotation of LP characters is directly linked to their legibility level. All LPs were manually labeled and then cross-validated using the PARSeq-tiny OCR model [19]. LPs classified as illegible — the lowest legibility level — are assigned an empty string, as their text cannot be recovered. For the remaining classes, the characters were annotated and validated through a semi-automatic process: the OCR model was used to predict each LP’s text, and any discrepancies between the manual and predicted annotations were reviewed and corrected.
@@
Table II presents the distribution of images in the dataset. While nighttime images constitute a minority of our dataset ( evening and nighttime images versus morning and afternoon images), this represents a significant contribution compared to other public datasets, which typically contain minimal nighttime imagery [20]. Most images contain at least one vehicle with a positive label (an LP that is legible, non-occluded, or associated with a valid vehicle), and images include at least one LP that is legible, valid, and not occluded.
@@ IV Experimental Setup
Our experiments focus on the legibility attribute at the LP level. We crop all LPs from the source images using their annotated corner coordinates. To evaluate legibility classification, we consider three models: ResNet [21], ViT [13] and YOLO-cls [14]. The goal is to assess whether an LP image is suitable for OCR processing, enabling more efficient downstream LP recognition by filtering out low-quality samples.
@@
We also assess the impact of SR on LP recognition performance. For SR, we adopt the pipelines proposed in [9], where images are first reconstructed using their SR model and then processed with a version of the GP_LPR OCR model [22] trained on the RodoSol-ALPR dataset [5]. Both SR and OCR are applied to all images, and the OCR results from the reconstructed images are compared against those from the originals. In this experiment, in addition to the LCOFL-GAN proposed in [9], we also evaluate Real-ESRGAN [15] and LPSRGAN [10], providing a broader overview of current SR state of the art.
Finally, we report OCR results by legibility class. In addition to GP_LPR, we also evaluate the PARSeq-tiny model [19], trained on an unpublished dataset. We report both character-level and LP-level recognition rate, presenting the results broken down by legibility class.
@@ IV-A Fine-Tuning Parameters
As mentioned before, we train ResNet, YOLO-cls, and ViT in these scenarios. Specifically, we employ ResNet-50 and ViT Base-16 from PyTorch’s torchvision library, and YOLO11m-cls from the ultralytics implementation. These models were selected based on their proven effectiveness and widespread adoption in state-of-the-art research across diverse domains [23, 24, 25]. We initialize all models with the weights pre-trained on ImageNet, and swap the last layer for a simple linear layer with outputs, where is the number of classes in a given training scenario (, , and classes for scenarios , , and , respectively).
@@ V Results
Table V presents the performance of the pre-trained SR models when integrated with LP recognition. All cropped LPs from the LPLC dataset were first enhanced using the respective SR model (LCOFL-GAN [9], RealESRGAN [15], and LPSRGAN [10]) and then processed with the GP_LPR OCR model [22], trained on the RodoSol-ALPR dataset [5] (as in [9]). This setup evaluates whether SR improves LP recognition accuracy. The results suggest otherwise: in the best case (Real-ESRGAN), recognition improved for only out of the LPs (). These findings indicate that current SR networks struggle to generalize across datasets.
@@
Finally, Table VI presents the character-wise accuracy (Char Acc) and whole-LP accuracy (LP Acc) for our dataset, broken down by legibility class. As described in Section IV, we use GP_LPR [22] and PARSeq-tiny [19], the latter of which was trained on a different, unpublished dataset. The results clearly indicate that OCR performance is strongly correlated with the legibility of the LP: lower legibility levels lead to significantly reduced recognition accuracy.
@@ Acknowledgments
This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) — Finance Code 001, and in part by the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) (# 315409/2023-1).

### v2

@@ Abstract
Automatic License Plate Recognition (ALPR) faces a major challenge when dealing with illegible license plates (LPs). While reconstruction methods such as super-resolution (SR) have emerged, the core issue of recognizing these low-quality LPs remains unresolved. To optimize model performance and computational efficiency, image pre-processing should be applied selectively to cases that require enhanced legibility. To support research in this area, we introduce a novel dataset comprising 10,210 images of vehicles with 12,687 annotated LPs for legibility classification (the LPLC dataset). The images span a wide range of vehicle types, lighting conditions, and camera/image quality levels. We adopt a fine-grained annotation strategy that includes vehicle- and LP-level occlusions, four legibility categories (perfect, good, poor, and illegible), and character labels for three categories (excluding illegible LPs). As a benchmark, we propose a classification task using three image recognition networks to determine whether an LP image is good enough, requires super-resolution, or is completely unrecoverable. The overall F1 score, below 80% for all three baseline models (ViT, ResNet, and YOLO), together with the analyses of SR and LP recognition methods, highlights the difficulty of the task and reinforces the need for further research. The proposed dataset is publicly available at https://github.com/lmlwojcik/lplc-dataset .
@@ I Introduction
Automatic License Plate Recognition (ALPR)plays a vital role in road surveillance by using automated systems to detect vehicles and recognize them based on their license plates (LPs) [laroca2021efficient, silva2022flexible]. In the deep learning era, key research tasks in this domain include LP detection and LP recognition, often complemented by additional processes such as vehicle detection and LP rectification [ke2023ultra, ding2024endtoend].
@@
State-of-the-art ALPR methods achieve over 95% accuracy in detection tasks and exceed 90% recognition rates under ideal conditions [laroca2022cross, liu2024improving]. However, their performance degrades significantly in challenging scenarios that are not well represented in most mainstream datasets. Such scenarios include low-light environments, adverse weather conditions (e.g., rain), and low-resolution or low-quality images caused by poor equipment or transmission compression artifacts [wahyu2024fog, nascimento2025toward].
@@
Furthermore, recent advances in super-resolution (SR) technology have also opened up new paths to deal with faulty images [nascimento2024enhancing, pan2024lpsrgan]. These models were adapted to an LP reconstruction task and were shown to aid in recovering information from otherwise illegible images. However, these models are often costly and not always needed, potentially even ruining otherwise suitable images for LP recognition (as we have observed in some of our experiments reported in this work).
@@
In this context, it is important to distinguish between image quality and character legibility. LP recognition performance depends not merely on image resolution or overall visual quality, but primarily on how legible the characters are. Fig. 1 illustrates this point: a visually high-quality image may contain illegible LPs due to factors such as camera distance, while a low-quality image may still feature clearly legible LPs. In some cases, a single image can include both legible and illegible LPs. We also observed that Optical Character Recognition (OCR) models often produce incorrect predictions with high confidence [laroca2023leveraging], particularly in cases where most characters are legible but one or two remain unclear (an issue also depicted in Fig. 1).
@@
With these challenges in mind, we present the License Plate Legibility Classification (LPLC) dataset111The LPLC dataset is available at https://github.com/lmlwojcik/lplc-dataset , which contains over 10k radar images collected from various locations across the Brazilian state of Paraná and includes more than 12k annotated LPs. The images exhibit a wide range of capture conditions, including varying lighting scenarios, vehicle types (motorcycles, cars, buses, and trucks), and image quality. A key contribution of this dataset is the assignment of each LP to one of four legibility levels: perfect, good, poor, or illegible. These labels provide a qualitative and inherently subjective assessment of LP legibility. For the first three categories, we also provide OCR annotations. The dataset includes both Brazilian and Mercosur LPs 222Following prior literature [silva2022flexible, laroca2023leveraging, nascimento2024enhancing], we use the term “Brazilian” to refer to the LP layout used in Brazil prior to the adoption of the Mercosur layout. and can be used as a benchmark for various ALPR-related tasks.
@@
In this work, we also use LPLC for a series of LP legibility assessment tasks using established image classification models: ResNet [prajwal2023comparative], ViT [dosovitskiy2021image] and YOLO-cls [ultralytics2025yolov11]. Additionally, we explore the use of SR by evaluating three state-of-the-art models [nascimento2024enhancing, wang2021realesrgan, pan2024lpsrgan] on our novel dataset. Building on these components, we design a recognition pipeline that incorporates a new decision step to determine whether an input image is sufficiently legible, requires further processing, or should be considered unrecoverable.
@@ II Related Work
State-of-the-art advancements in ALPR, as in other machine learning research domains, are quantitatively assessed using public benchmark datasets [laroca2022cross, ke2023ultra]. An example is CLPD [clpd], a dataset containing 1,200 images captured with various devices in mainland China. While it primarily features daylight scenes and passenger cars, it offers considerable variability in backgrounds, road types, capture devices, and image quality.
@@
For the Brazilian context, the RodoSol-ALPR dataset [laroca2022cross] comprises images captured at toll plazas, providing a realistic representation of actual operating conditions. It includes a wide range of lighting scenarios, vehicle types, and weather conditions. The authors [laroca2022cross] demonstrated that state-of-the-art methods perform well on older datasets featuring simpler scenarios, such as Caltech-Cars [caltech], but struggle on RodoSol-ALPR, revealing limitations tied to outdated benchmarks. While RodoSol-ALPR captures real situations, it is important to note that these toll plaza environments differ significantly from many surveillance camera settings, where camera quality and capture conditions are often much poorer and more variable.
@@
Recent advances in LP recognition have shifted focus from achieving high performance on simpler datasets to addressing more specific challenges in the field. For example, Liu et al. [liu2024improving] proposed an attention-based decoder for LP recognition, following the growing trend of integrating attention mechanisms. Their approach combines a CNN encoder for feature extraction with a contrastive learning strategy designed to differentiate both the position and class of each character. The use of attention is motivated by the fixed character positions in LPs of a given format, enabling more effective layout-aware feature learning. They reported a % recognition rate on RodoSol-ALPR.
@@
Similarly, Rao et al. [crnn] proposed an end-to-end pipeline that includes a segmentation step. The method begins with a YOLOv5-based LP detector, followed by segmentation using the proposed AFF-Net model. Next, a skew correction module is applied to the segmentation map using standard image processing techniques to extract the LP region. The four corners of this region are then aligned with those of a canonical LP format. Finally, LP recognition is performed using a CNN. This correction step is effective in handling severely skewed LPs captured from oblique angles, such as those found in the CLPD dataset, where they reported a recognition rate of .
@@
SR has gained interest in ALPR research due to its potential to recover otherwise illegible LP images. For example, studies like [pan2024lpsrgan] adapt image-to-image translation methods, such as Generative Adversarial Network (GAN), to improve recognition performance on low-resolution LPs.
@@
In the same direction, Nascimento et al. [nascimento2024enhancing] proposed a novel focal loss tailored for character reconstruction, trained on a variant of the RodoSol-ALPR dataset called RodoSol-SR. This dataset comprises pairs of high-resolution LP images and their synthetically degraded low-resolution counterparts. These pairs were generated by the authors to train the model to reverse the degradation process through reconstruction.
@@ III The LPLC Dataset
We introduce the License Plate Legibility Classification (LPLC) dataset, a publicly available resource primarily designed for the task of classifying the legibility of LP images, that is, whether an LP image is suitable for direct OCR or requires additional processing. Although its main purpose is legibility classification, the dataset is also suitable for other tasks such as LP detection and recognition, due to its fine-grained annotations that include details on LP text, legibility levels, and occlusion. All annotations were made manually by a single person (the first author) using the VIA annotation software [dutta2019vgg], and validated through a semi-automatic process to ensure accuracy. This revision process involved using an OCR model’s outputs (PARSeq-tiny [bautista2022scene]) and identifying legibility classification errors to flag potentially inconsistent annotations in both the legibility labels and OCR transcripts. Fig. 2 presents a few sample images from the dataset.
@@
The annotation of LP characters is directly linked to their legibility level. All LPs were manually labeled and then cross-validated using the PARSeq-tiny OCR model [bautista2022scene]. LPs classified as illegible — the lowest legibility level — are assigned an empty string, as their text cannot be recovered. For the remaining classes, the characters were annotated and validated through a semi-automatic process: the OCR model was used to predict each LP’s text, and any discrepancies between the manual and predicted annotations were reviewed and corrected.
@@
Table II presents the distribution of images in the dataset. While nighttime images constitute a minority of our dataset ( evening and nighttime images versus morning and afternoon images), this represents a significant contribution compared to other public datasets, which typically contain minimal nighttime imagery [onim2023unleashing]. Most images contain at least one vehicle with a positive label (an LP that is legible, non-occluded, or associated with a valid vehicle), and images include at least one LP that is legible, valid, and not occluded.
@@ IV Experimental Setup
Our experiments focus on the legibility attribute at the LP level. We crop all LPs from the source images using their annotated corner coordinates. To evaluate legibility classification, we consider three models: ResNet [he2016deep], ViT [dosovitskiy2021image] and YOLO-cls [ultralytics2025yolov11]. The goal is to assess whether an LP image is suitable for OCR processing, enabling more efficient downstream LP recognition by filtering out low-quality samples.
@@
We also assess the impact of SR on LP recognition performance. For SR, we adopt the pipelines proposed in [nascimento2024enhancing], where images are first reconstructed using their SR model and then processed with a version of the GP_LPR OCR model [liu2024irregular] trained on the RodoSol-ALPR dataset [laroca2022cross]. Both SR and OCR are applied to all images, and the OCR results from the reconstructed images are compared against those from the originals. In this experiment, in addition to the LCOFL-GAN proposed in [nascimento2024enhancing], we also evaluate Real-ESRGAN [wang2021realesrgan] and LPSRGAN [pan2024lpsrgan], providing a broader overview of current SR state of the art.
Finally, we report OCR results by legibility class. In addition to GP_LPR, we also evaluate the PARSeq-tiny model [bautista2022scene], trained on an unpublished dataset. We report both character-level and LP-level recognition rate, presenting the results broken down by legibility class.
@@ IV-A Fine-Tuning Parameters
As mentioned before, we train ResNet, YOLO-cls, and ViT in these scenarios. Specifically, we employ ResNet-50 and ViT Base-16 from PyTorch’s torchvision library, and YOLO11m-cls from the ultralytics implementation. These models were selected based on their proven effectiveness and widespread adoption in state-of-the-art research across diverse domains [yolo_app, lima2024toward, laroca2025improving]. We initialize all models with the weights pre-trained on ImageNet, and swap the last layer for a simple linear layer with outputs, where is the number of classes in a given training scenario (, , and classes for scenarios , , and , respectively).
@@ V Results
Table V presents the performance of the pre-trained SR models when integrated with LP recognition. All cropped LPs from the LPLC dataset were first enhanced using the respective SR model (LCOFL-GAN [nascimento2024enhancing], RealESRGAN [wang2021realesrgan], and LPSRGAN [pan2024lpsrgan]) and then processed with the GP_LPR OCR model [liu2024irregular], trained on the RodoSol-ALPR dataset [laroca2022cross] (as in [nascimento2024enhancing]). This setup evaluates whether SR improves LP recognition accuracy. The results suggest otherwise: in the best case (Real-ESRGAN), recognition improved for only out of the LPs (). These findings indicate that current SR networks struggle to generalize across datasets.
| GAN Model |
Class | Total | ||||
| Perfect | Good | Poor | |||||
| LCOFL-GAN [nascimento2024enhancing] | Better | ||||||
| Equal | |||||||
| Worse | |||||||
| Real-ESRGAN [wang2021realesrgan] | Better | ||||||
| Equal | |||||||
| Worse | |||||||
| LPSRGAN [pan2024lpsrgan] | Better | ||||||
| Equal | |||||||
| Worse | |||||||
@@
Finally, Table VI presents the character-wise accuracy (Char Acc) and whole-LP accuracy (LP Acc) for our dataset, broken down by legibility class. As described in Section IV, we use GP_LPR [liu2024irregular] and PARSeq-tiny [bautista2022scene], the latter of which was trained on a different, unpublished dataset. The results clearly indicate that OCR performance is strongly correlated with the legibility of the LP: lower legibility levels lead to significantly reduced recognition accuracy.
| Class | GP_LPR [liu2024irregular] | PARSeq-tiny [bautista2022scene] | ||
| Char Acc | LP Acc | Char Acc | LP Acc | |
| Perfect | % | % | % | % |
| Good | % | % | % | % |
| Poor | % | % | % | % |
| Overall | % | % | % | % |
@@ Acknowledgments
This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES), through the Programa de Excelência Acadêmica (PROEX) - Finance Code 001, and in part by the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) (# 315409/2023-1). The authors gratefully thank these organizations for the support.